OpenStudy (anonymous):

what does the graph of -3x-5y is greater than or equal to 20 look like? please help.

OpenStudy (anonymous):

\[-3x-5y \ge20\]

zepdrix (zepdrix):

\[\Large\rm -3x-5y\ge 20\]Ummm I always find these easier to work with if you put them in slope-intercept form. Add 3x to each side,\[\Large\rm -5y\ge 3x+20\]Divide each side by -5,\[\Large\rm y\le -\frac{3}{5}x-4\] y-intercept of -4, negative slope.

zepdrix (zepdrix):

So something like this,

zepdrix (zepdrix):

So let's deal with the inequality so we can get some shading and all that.

zepdrix (zepdrix):

Oh oh back up a sec, do you understand why I FLIPPED the inequality sign? We divided a negative across the inequality. We have the `or equal to` included in our inequality, so do we want a `solid` or `dashed` line?

OpenStudy (anonymous):

Yes I understand why you flipped the inequality. What I am having trouble knowing is what makes the line be solid or dashed. BTW, thanks for helping me.

zepdrix (zepdrix):

\[\Large\rm y=-\frac{3}{5}x-4\]^This would be the line. The inequality is also giving us shading on one of the sides. See how the inequality has the `or equal to` on it? That means we want to `include` the values along the line. So it should be a SOLID line, we're including all those values.